OpenStudy (anonymous):

What are the zeros of the function? Graph the function. A. –3, 3, –4 B. –3, 3, 4 C. 3, –3, –4 D. 3, –3, -4

ganeshie8 (ganeshie8):

where is the function ?

OpenStudy (anonymous):

y=(x+3)(x-3)(x-4)

ganeshie8 (ganeshie8):

set y=0 to find the zeroes : 0 = (x+3)(x-3)(x-4)

ganeshie8 (ganeshie8):

at least one of those factors must equal 0 for the entire product to be 0 : 0 = x+3 or 0 = x-3 or 0 = x-4